OpenStudy (anonymous):

plz help me with an 8th grade question

OpenStudy (anonymous):

The steps below show the incomplete solution to find the value of x for the equation 7x + 8 − 3x = −6 + 10: Step 1: 7x + 8 − 3x = −6 + 10 Step 2: 7x + 8 − 3x = 4 Step 3: 4x + 8 = 4 Which of these is most likely the next step? 4x = 12 4x = 4 4x = −4 4x = −12

OpenStudy (anonymous):

@InExileWeTrust @AaronAndyson @Danni_joanaveil @Quiroz15Army @tylermcmullen23 @nechirwan

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

any guesses?

OpenStudy (anonymous):

You now need to take the 8 from both sides: 4x + 8 - 8 = 4 - 8 Which equals: 4x = -4

OpenStudy (kidrah69):

Combine your like terms we want to simplyfy it so subtract 8 from both sides

OpenStudy (anonymous):

also i need help with this one 2 What is the value of y in the equation 2(4y − 1) = 6?

OpenStudy (kidrah69):

distribute 2 x 4 =? 2 x -1 =?

OpenStudy (anonymous):

nvm it is 1

OpenStudy (anonymous):

thank u

OpenStudy (skullpatrol):

do know how to check it?